A company gives each of its 50 employees (assume they were all employed continuously through 2017 and 2018) 12 days of vacation a year if they are employed at the end of the year. The vacation accumulates and may be taken starting January 1 of the next year. The employees work 8 hours per day. In 2017, they made $36 per hour and in 2018 they made $39 per hour. During 2018, they took an average of 9 days of vacation each. The company’s policy is to record the liability existing at the end of each year at the wage rate for that year. What amount of vacation liability would be reflected on the 2017 and 2018 balance sheets, respectively?
Answer #1

DETERMINATION OF AMOUNT OF VACATION LIABILITY REFLECTED ON THE 2017 AND 2018 BALANCE SHEETS.

Vacation liability

= NO. of employees × Vacation days × working hours per day × hourly wage rate.

For 2017 :

= 50 × 12 × 8 × $36 = $172800.

For 2018

= 50 × 9 × 8 × $39 = $140400.